write an equation in slope-intercept form for the line with slope -4 that contains the point (0, -3)
Answer by Cromlix(4381)

Hi
Using formula: y - b = m(x - a) m = -4 and (a, b) = (0, -3)
y - (-3) =-4(x - 0)
y + 3 = -4x
y = -4x - 3
Hope this helps :-)